2 charged particles reacted with a force of 0.31 N with a distance of 1.5 cm, the 2nd charge contains 54 nC. Find the value of the first charge


1
Expert's answer
2022-03-08T12:31:14-0500

Answer

First charge is

"Q=\frac{Fr^2}{kQ'}"

"Q=\frac{0.31*0.015*0.015}{9*10^9*54*10^{-9}}\\=0.14\mu C"